OpenStudy (anonymous):

find all answers for cos(tetha)+1=0

OpenStudy (anonymous):

\[\theta=\pi, 3 \pi\]

OpenStudy (anonymous):

steps ?

OpenStudy (anonymous):

cos\[\cos \theta=-1\]

OpenStudy (anonymous):

ik did that but arent you supposed to do cos inverse of 1 not -1 ???

OpenStudy (anonymous):

\[\theta =\cos^{-1}( -1)\]

OpenStudy (anonymous):

no u can

OpenStudy (anonymous):

u can totally understand if u see the cos graph

OpenStudy (anonymous):

thanks :D i got it but shouldnt be 2pi ?

OpenStudy (anonymous):

no 2pi -pi=pi...nxt one is 2pi+pi=pi

OpenStudy (anonymous):

u will get pi 2 tyms,bt u hv to mention it one tym

OpenStudy (anonymous):

ok thanks